# Retrieval and Distill: A Temporal Data Shift-Free Paradigm for Online Recommendation System

### v4

@@ 4.3.2. Sequential based Models
Table 2. Performance comparison of CTR prediction task. Rel.Impr means relative AUC improvement rate against each original model. Improvements are statistically significant with $p<0.01$.
@@
relevance network pretrained with $\mathcal{D}_{\text{train}}$ DNN 0.6210 0.7139 0.3373 0.9291 0.6227 0.7070
@@
relevance network pretrained with $\mathcal{D}_{\text{shifting}}$ DNN 0.5224 0.8163 0.3141 0.9398 0.5006 0.8396
@@
Table 3. Performance comparison of pretrained relevance network in Retrieval Framework.
Loss Functions Taobao Tmall Alipay
Mean Squared Error loss DNN 0.7224 0.8961 0.7423
DeepFM 0.7192 0.8961 0.7532
PNN 0.7269 0.9026 0.7713
Kullback-Leibler loss DNN 0.7238 0.8937 0.7722
DeepFM 0.7220 0.8940 0.7661
PNN 0.7277 0.9013 0.7865
Cosine Similarity loss DNN 0.7274 0.8936 0.7649
DeepFM 0.7248 0.8940 0.7593
PNN 0.7301 0.9030 0.7791
Table 4. Distill Framework AUC of different distillation loss functions.
4.4. Overall Performance (RQ1)
@@
4.5. Applying Knowledge Distillation to Non-Parametric Models (RQ2)
@@
4.6. Temporal Invariance of Association (RQ3)
@@
4.7. Performance Efficiency Study (RQ4)
@@
Recently, a range of retrieval-based models have been introduced in the field of recommender systems, some of which utilize discrete features for retrieval, such as SIM (Pi et al., 2019) and RIM(Qin et al., 2021), while others employ continuous features, like DERT(Zheng et al., 2023). Due to the significant time overhead associated with retrieval, these models face challenges in being directly deployed online. In Figure 6, we compare three methods for augmenting relevant information. As RAD employs distillation to bypass the retrieval process, it incurs no retrieval time. On the other hand, directly aggregating the search-distill module with the original model increases the inference burden on the original model. However, as shown in Figure 7, the additional inference load introduced by incorporating the search-distill module is not significant, and the time overhead it introduces is considerably less than the burden added by using sequence models for model inference.
4.8. The Study of Distill Loss Functions (RQ5)
In Table 4, we compare the impact of various distillation functions on the Distill Framework. For simplicity, our experiments were conducted using only three models based on feature interaction. Our findings indicate that the performance enhancement provided by different distillation loss functions is broadly similar, suggesting that the quantity of information extracted primarily depends on the shifting data itself. However, slight variations exist in the performance of different loss functions across datasets. The Mean Squared Error Loss Function generally performs slightly better on Tmall, the Kullback-Leibler Divergence Loss Function shows slight superiority on Alipay, and the cosine similarity loss function performs marginally better on Taobao. When addressing different data distributions, selecting an appropriate distillation function tailored to the specific dataset can further enhance the distillation effect.
Figure 6. Comparison of Retrieval time.
Figure 7. Comparison of Inference time.
5. Related Work
@@
Click-through rate (CTR) prediction is crucial in the fields of online advertising, recommendation systems, and information retrieval. Many deep learning models for CTR prediction primarily concentrate on investigating the feature interactions within a single instance (Guo et al., 2017; He et al., 2017; Qu et al., 2018; Liu et al., 2020; Gorishniy et al., 2021). The Wide&Deep network (Cheng et al., 2016) constructs fully connected (FC) layers based on field feature embeddings to extract feature interactions implicitly. CCPM (Liu et al., 2015) employs convolution operators to probe local feature interaction patterns from a list of feature embeddings. PNN (Qu et al., 2018) suggests executing inner or outer product operations on each pair of field feature embeddings, followed by FC layers for further implicit feature interaction extraction. DeepFM (Guo et al., 2017) evolves from the factorization machine (FM) (Rendle, 2010). Recently, graph neural networks have also shown great potential in recommendation scenes by constructing data into graph structures and performing message passing (Li et al., 2019; Wang et al., 2019b; He et al., 2020; Wang et al., 2019a; Wu et al., 2021). Despite the numerous models proposed in this direction, the model capacity might be constrained as the model’s input is a single instance, necessitating the model to encode all knowledge into the parameters. Our work employs these deep models as the base model of the framework.
@@
Retrieval-based methods for CTR prediction are typically derived from sequential recommendation tasks. where it’s common to use a recurrent neural network (RNN) to construct user profile representations for recommendation predictions. However, when the user’s historical behavior sequence becomes significantly long, RNN architectures struggle to memorize and utilize early behavior patterns (Pi et al., 2019). Therefore, retrieval methods serve as a practical alternative to lengthy recurrence and attention operations. UBR (Qin et al., 2020) creates a search index from the user’s historical sequence data and employs reinforcement learning to generate a search query based on current prediction data. The retrieved data is then aggregated through an attention operation and input into an MLP for label prediction. SIM (Qi et al., 2020) executes both hard search based on item category information and soft search based on item embedding vectors, using the DIEN model (Zhou et al., 2019) as the backbone. RIM (Qin et al., 2021) expands the research scenario from sequential recommendation to generic tabular data prediction. The query is the target row data, and the retrieved data instances are the relevant rows in the table, with BM25 (Robertson et al., 1995) serving as the predefined relevance function between two rows. DERT (Zheng et al., 2023) is the pioneering work to incorporate dense retrieval into recommendation systems. Retrieval-enhanced machine learning is now considered a significant frontier in machine learning research (Zamani et al., 2022).
@@
Knowledge Distillation (KD) (Hinton et al., 2015) seeks to transfer the ’dark knowledge’ from a teacher model to a student model and is extensively used in model compression and knowledge transfer (Gou et al., 2021). Recent studies have also employed KD in various recommendation tasks, such as click-through rate (CTR) prediction (Zhu et al., 2020), topology distillation (Kang et al., 2021), and multi-task transfer (Yang et al., 2022a). Prior theoretical analyses (Tang et al., 2020; Phuong and Lampert, 2019; Chandrasegaran et al., 2022) have indicated that the teacher’s predictions, also referred to as soft labels, are typically viewed as more informative training signals compared to binary hard labels. KD can adaptively adjust the training dynamics of the student model based on the values of the soft labels. Knowledge distillation is typically employed to transfer knowledge from a larger parameterized model to a smaller one. Recent studies have applied knowledge distillation to transfer knowledge from KNN to parameterized models (Zhang et al., 2023; Yang et al., 2022b), demonstrating the potential of knowledge distillation in non-parametric models.
6. Conclusion
@@
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v5

@@ 4.3 Compared Methods
Table 1: Performance comparison of CTR prediction task. Rel.Impr means relative AUC improvement rate against each original model. Improvements are statistically significant with $p<0.01$.
@@
pretrained with $\mathcal{D}_{\text{train}}$ DNN 0.6210 0.7139 0.3373 0.9291 0.6227 0.7070
@@
pretrained with $\mathcal{D}_{\text{shifting}}$ DNN 0.5224 0.8163 0.3141 0.9398 0.5006 0.8396
@@
Table 2: Performance comparison of pretrained relevance network in Retrieval Framework.
4.4 Overall Performance (RQ1)
@@
4.5 Applying Knowledge Distillation to Non-Parametric Models (RQ2)
@@
4.6 Temporal Invariance of Association (RQ3)
@@
4.7 Performance Efficiency Study (RQ4)
@@
Recently, a range of retrieval-based models have been introduced in the field of recommender systems, some of which utilize discrete features for retrieval, such as SIM [21] and RIM[24], while others employ continuous features, like DERT[44]. Due to the significant time overhead associated with retrieval, these models face challenges in being directly deployed online. In Figure 6, we compare three methods for augmenting relevant information. As RAD employs distillation to bypass the retrieval process, it incurs no retrieval time. On the other hand, directly aggregating the search-distill module with the original model increases the inference burden on the original model. However, as shown in Figure 7, the additional inference load introduced by incorporating the search-distill module is not significant, and the time overhead it introduces is considerably less than the burden added by using sequence models for model inference.
Figure 6: Comparison of Retrieval time.
Figure 7: Comparison of Inference time.
5 Related Work
@@
Click-through rate (CTR) prediction is crucial in the fields of online advertising, recommendation systems, and information retrieval. Many deep learning models for CTR prediction primarily concentrate on investigating the feature interactions within a single instance [6, 8, 25, 17]. The Wide&Deep network [2] constructs fully connected (FC) layers based on field feature embeddings to extract feature interactions implicitly. Recently, graph neural networks have also shown great potential in recommendation scenes by constructing data into graph structures and performing message passing [16][36][7][35][37]. Despite the numerous models proposed in this direction, the model capacity might be constrained as the model’s input is a single instance, necessitating the model to encode all knowledge into the parameters. Our work employs these deep models as the base model of the framework.
@@
Retrieval-based CTR prediction methods, adapted from sequential recommendation, often replace RNNs when processing long user behavior sequences, as RNNs struggle with early pattern recall [21], offering a more efficient alternative to recurrence and attention mechanisms. RIM [24] expands the research scenario from sequential recommendation to generic tabular data prediction. The query is the target row data, and the retrieved data instances are the relevant rows in the table, with BM25 [27] serving as the predefined relevance function between two rows. DERT [44] is the pioneering work to incorporate dense retrieval into recommendation systems. Retrieval-enhanced machine learning is now considered a significant frontier in machine learning research [42].
@@
Knowledge Distillation (KD) [11] seeks to transfer the ’dark knowledge’ from a teacher model to a student model and is extensively used in model compression and knowledge transfer [4]. Recent studies have also employed KD in various recommendation tasks, such as click-through rate (CTR) prediction [48], topology distillation [13], and multi-task transfer [38]. Prior theoretical analyses [30, 20, 1] have indicated that the teacher’s predictions, also referred to as soft labels, are typically viewed as more informative training signals compared to binary hard labels. KD can adaptively adjust the training dynamics of the student model based on the values of the soft labels. Knowledge distillation is typically employed to transfer knowledge from a larger parameterized model to a smaller one. Recent studies have applied knowledge distillation to transfer knowledge from KNN to parameterized models [43, 39], demonstrating the potential of knowledge distillation in non-parametric models.
6 Conclusion